OpenStudy (anonymous):

Add or Subtract. 5 (square root of 3) + square root of 48

OpenStudy (anonymous):

@agent0smith D oyou know how to do this?

OpenStudy (anonymous):

\[5\sqrt{3} + \sqrt{48}\]

OpenStudy (agent0smith):

Simplify the square root of 48 first... 48 = 16*3 sqrt(16*3) can be simplified

OpenStudy (anonymous):

I got \[5\sqrt{3}+4\sqrt{3}\]

OpenStudy (anonymous):

So I would add up 5 and 4 giving me 9 radical 3?

OpenStudy (agent0smith):

Yep.

OpenStudy (anonymous):

Thankyou! Can you help me with a couple more please?

OpenStudy (agent0smith):

If i can't, someone will :)

OpenStudy (anonymous):

haha true!